2. (CMA) Given the following information for Xerbert Company (in thousands): Static Budget for 2008 Actual Results for 2008 Xenox Xeon Total Xenox Xeon Total 150 100 250 130130 260 Units sold Revenues Variable costs Contribution margin S450 S 250 S 700 S390 S 260 S 650 $900 $1,000 $1,900 $780 $1,235 $2,015 450 750 1,200 390 975 1365 Fixed costs: Manufacturing Marketing Customer service Total fixed costs Operating income 200 153 95 448 S 252 190 140 90 420 S 230 a. b. c. Compute the sales-volume variance for both products together. Compute the sales-mix variance for both products together. Compute the sales-quantity variance for both products together.
Answer #1
Static Budget for 2018 Actaul results for 2018
Xenox Xeon Total Xenox Xeon Total
Units sold 150 100 250 130 130 260
Revenues ($) 900 1000 1900 780 1235 2015
Variable Cost ($) 450 750 1200 390 975 1365
Contribution ($) 450 250 700 390 260 650
Selling Price per unit 6 10 7.6 6 9.5 7.75
Contribution per unit 3 2.5 2.8 3 2 2.5

a. Sales Volume Variance represents the difference between the actual units sold and the budgeted quantity multiplied by either the standard profit per unit or standard contribution per unit. In Absorption costing standard profit per unit is used, but in marginal costing, standard contribution per unit must be used.

Sales Volume Variance = Standard Contribution per unit * (Standard Quantity - Actual Quantity)

For Xenox = 3 *(150-130) = 60 Adverse
For Xeon = 2.5*(100-130) = 75 Favourable
Combined =60 Adverse + 75 Favourable = 15 Favourable

b. Sales Mix Variance arise when the company manufactures and sells more than one type of product. This variance will be due to actual mix and budgeted mix of sales.

  Sales Mix Variance = Standard Contribution per unit * (Standard proportion of actual sales - Budget quantity of sales)

For Xenox = 3*[(150/250*260) - 130] = 78 Favourable
For Xeon = 2.5*[(100/250*260)-130] = 65 Adverse
Combined = 78 Adverse + 65 Favourable = 13 Favourable


c. Sales Quantity Variance is the difference between budgeted profit on budgeted sales and expected profit on actual sales.

Sales Quantity Variance =  Standard Contribution per unit * ( Revised Standard Quantity - Standard Quantity)

For Xenox = 3*[(150/250*260]-150 = 18 Favourable
For Xeon = 2.5*[(100/250*260)-100] = 10 Favourable
Combined = 18 Favourable + 10 Favourable = 28 Favourable